Jack and Jill's age has a total of 62. In 2 years, Jack will be 9 years less than twice Jill's age. How old are they now?
Answer by stanbon(75887)

Jack and Jill's age has a total of 62. In 2 years, Jack will be 9 years less than twice Jill's age. How old are they now?
----------------
Equations:
J + j = 62
J+2 = 2(j+2)-9
--------------------
Substitutute for "J" and solve for "j"::
(62-j)+2 = 2j+4+9
3j = 64 - 13
3j = 51
j = 17 yrs (Jill's age now)
---
Solve for "J":
J + j = 62
J = 62-17 = 45 years
========================
Cheers,
Stan H.
-----------------